For each of the following solutions, calculate the:

A.) grams of 1.6 % (m/m) NaCl solution that contains 6.00 g of NaCl

Express your answer to two significant figures and include the appropriate units.

B.) milliliters of 45 % (m/v) NaF solution that contains 6.9 g of NaF

Express your answer to two significant figures and include the appropriate units.

C.) milliliters of 8.9 % (v/v) ethanol solution that contains 15.7 mL of ethanol

Express your answer to two significant figures and include the appropriate units.
Answer #1

A)

(m/m)% = mass of NaCl * 100 / mass of solution

1.6 = 6.00 * 100 / mass of solution

mass of solution = 375 g

Answer: 3.7*10^2 g

B)

(m/v)% = mass of NaF * 100 / volume of solution

45 = 6.9 * 100 / volume of solution

mass of solution = 15 mL

Answer: 15 mL

C)

(v/v)% = volume of ethanol * 100 / volume of solution

8.9 = 15.7 * 100 / volume of solution

mass of solution = 176 mL

Answer: 1.8*10^2 mL